OpenStudy (anonymous):

If f(x)=x^3 , simplify the expression f(x+h)-f(x)/h

OpenStudy (anonymous):

\[\Large \lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

Can you plug in the numbers?

geerky42 (geerky42):

He only need to simplify expression, not finding limit.

OpenStudy (anonymous):

He doesn't need the derivative? So what exactly does he want to do? Linear approximation at some x=a? or what? looks like he needs the derivative.

geerky42 (geerky42):

Simplify the expression...?

OpenStudy (tkhunny):

#1 - You must learn your Order of Operations f(x+h)-f(x)/h MEANS \(f(x+h) - \dfrac{f(x)}{h}\) and you will have a tough time simplifying that. #2 - Substitute the expression and simplify it. [f(x+h)-f(x)]/h MEANS \(\dfrac{f(x+h) + f(x)}{h}\) and you will have ta chance to succeed at the exercise. You have \(f(x) = x^{3}\) - First substitute!

OpenStudy (anonymous):

Yes just simplify.

geerky42 (geerky42):

seriously, @HELP_ME! ?

geerky42 (geerky42):

ok I help you. What do you have so far?

OpenStudy (anonymous):

@geerky42 I need your help

OpenStudy (anonymous):

Well all of you have been slightly confusing so I dont have anything as of now.

geerky42 (geerky42):

Ok, refresh your brain, forgot what you read. You are given that \(f(x) = x^3\), right? And you were asked to simplify \(\dfrac{f(x+h)-f(x)}{h}\). Because \(f(x)=x^3\), you have \(f(x+h) = (x+h)^3\) which is \( (x+h) (x+h) (x+h)\) Now substitute and simplify.

OpenStudy (tkhunny):

Since you decided to make the warning public, I choose to defend it publicly. It was just a warning. Try to learn from it and be angry somewhere else. There IS no need to argue with the question or to swear at the teachers. Can we get back to learning mathematics?

OpenStudy (anonymous):

Okay I am sorry

OpenStudy (anonymous):

So now we have (x+h)(x+h)(x+h) ------------ h so we can cross cancel one of the h's and that leaves us (x)(x+h)(x+h) ------------ 1 which we can turn into (x)(x+h)^2 ???? did I do it?

geerky42 (geerky42):

Well, you should have \(\dfrac{ (x+h) (x+h) (x+h)-x^3}{h}\) First, try to expand \( (x+h) (x+h) (x+h)\)

OpenStudy (anonymous):

Follow these steps in which I used that that \[ (a + b)^3 a^3 + 3 a^2 b + 3 a b^2 + b^3 \] \[\frac{f(h+x)-f(x)}{h}=\frac{(h+x)^3-x^3}{h}=\\ \frac{h^3+3 h^2 x+3 h x^2+x^3-x^3}{h}=\\ \frac{h^3+3 h^2 x+3 h x^2+x^3-x^3}{h}\\=\frac{h \left(h^2+3 h x+3 x^2\right)}{h}=h^2+3 h x+3 x^2\]

OpenStudy (anonymous):

Typo \[ (a + b)^3 = a^3 + 3 a^2 b + 3 a b^2 + b^3 \]

OpenStudy (anonymous):

This limit above will help you prove that the derivative of \(x^3\) is \( 3 x^2 \)

OpenStudy (anonymous):

I am only trying to simplify though.

OpenStudy (anonymous):

Read my post above the answer is \[ h^2 + 3 h x + 3 x^2 \]

OpenStudy (anonymous):

Are you positive? and could I get a second opinion from someone? I am just confused.

OpenStudy (anonymous):

I am positive 100%

geerky42 (geerky42):

I cam confirm that answer is \(3x^2+3xh+h^2\)

geerky42 (geerky42):

What part are you confused with? we can explain more

OpenStudy (anonymous):

You should try to redo all the steps in my answer on your own to learn this stuff

OpenStudy (anonymous):

Okay I thank you both, but I was simply confused on combining the x and the h but I guess that you foiled?

geerky42 (geerky42):

basically, yeah. \(=~(x+h)x^2+(x+h)2xh+(x+h)h^2\) Simplify further and combine like terms.